OpenStudy (anonymous):

Find a and b such that y=(x+a)/(x+b)^2 has a local minimum/maximum at x=M and a point of inflection at x=N. (I need hints please)

OpenStudy (anonymous):

1/ take y' and let y'=0 , replace x =M into the expression of y'=0 (you may have 1 equation) 2/ take y" and let y"=0, replace x =N, you may have the 2nd equation 2equations, 2 unknowns a and b. --> it's solvable

OpenStudy (anonymous):

\[y' = \frac{ (x+b)^2 - (2)(x+a)(x+b) }{ (x+b)^4 }\]

OpenStudy (anonymous):

my mind tells me I can factor out in there but my sleepy sleepy mind is in doubt and researching

OpenStudy (anonymous):

\[y'=\dfrac{-2a+b-x}{(b+x)^3}\] no condition since we have x = M not -b, --> y'=0 iff -2a+b -M =0 (this is the first equation) \[y"=\dfrac{2(3a-2b+x)}{(b+x)^4}\] so that y"=0 iff 3a-2b+N=0 (this is the second one) solve for a, b from those equations

OpenStudy (anonymous):

show me your work, please. The answers should be respect to M, N

OpenStudy (anonymous):

I got a= N - 2M and b= 2N - 3M

OpenStudy (anonymous):

I just solved a twice verifying and got a = N - 2M

OpenStudy (anonymous):

Oh, I am sorry, you are correct